OpenStudy (anonymous):

Determine the next term of the sequence and determine if the sequence is arithmetric, geometric, or neither. 3,18,108,648,3888 The next term is ...? A. Arithmetric B. Geometric C. Neither arithmetic nor geometric

jimthompson5910 (jim_thompson5910):

If it was arithmetic, then the difference between each consecutive term would be the same. However, notice that 18-3 = 15 and 108 - 18 = 90 These differences are not equal. So it's not an arithmetic sequence. ------------------- If it was a geometric sequence, then the ratio of each term with its previous term would be the same. So divide each term by its previous term: 18/3 = 6 108/18 = 6 648/108 = 6 3888/648 = 6 This points us to the fact that this is indeed a geometric sequence. The next term would then be (Common Ratio)*(Last term) = 6*3888 = 23328

OpenStudy (anonymous):

Thanks Jim

jimthompson5910 (jim_thompson5910):

no problem